OpenStudy (anonymous):

Express the confidence intervals (0.026, 0.124) in the form of p - E < p < p + E.

OpenStudy (dan815):

what is P and E?

OpenStudy (anonymous):

\[\frac{(p-E)+(p+E)}{2}=p\] where \(p\) is the midpoint of a confidence interval. You're explicitly given that \(p-E=0.026\) and \(p+E=0.124\). You can easily find \(p\) then \(E\) from here. Oh, and the interval should be in the form \(p-E<\hat{p}<p+E\), where \(\hat{p}\) is whatever statistic is being estimated.